### v3

# Safe Navigation in Uncertain Crowded Environments Using Risk Adaptive CVaR Barrier Functions

Abstract
Robot navigation in dynamic, crowded environments poses a significant challenge due to the inherent uncertainties in the obstacle model. In this work, we propose a risk-adaptive approach based on the Conditional Value-at-Risk Barrier Function (CVaR-BF), where the risk level is automatically adjusted to accept the minimum necessary risk, achieving a good performance in terms of safety and optimization feasibility under uncertainty. Additionally, we introduce a dynamic zone-based barrier function which characterizes the collision likelihood by evaluating the relative state between the robot and the obstacle. By integrating risk adaptation with this new function, our approach adaptively expands the safety margin, enabling the robot to proactively avoid obstacles in highly dynamic environments. Comparisons and ablation studies demonstrate that our method outperforms existing social navigation approaches, and validate the effectiveness of our proposed framework. [Paper Page][Video][Code].
I Introduction
Safe navigation in crowded environments with dynamic obstacles remains a fundamental challenge in robotics due to obstacle uncertainty. The common approaches involve using risk metrics to quantify and enforce safety constraints under this uncertainty, such as conventional stochastic control methods [1, 2] and robust control methods [3, 4]. While robust control methods prioritize worst-case scenarios, often leading to overly conservative behaviors, stochastic methods optimize for expected performance, which may result in unsafe decisions in high-risk situations.
Recently, [5, 6, 4, 7] have explored combining Conditional Value-at-Risk (CVaR) with Control Barrier Functions (CBFs) to achieve probabilistic safety, thereby providing a trade-off between conservatism and efficiency. In their frameworks, CBFs [8] act as safety filters that regulate control inputs, while CVaR [9] quantifies the expected risk under a given risk level. However, existing methods based on the CVaR barrier function (CVaR-BF) primarily rely on a fixed risk level that remains constant throughout the trajectory. This can limit their applicability in dynamic environments with crowded obstacles. We highlight that such a fixed risk level is not flexible enough: A low risk tolerance enhances safety but can render the optimization infeasible, whereas a high risk tolerance improves feasibility at the expense of safety.
To address this limitation, we propose a risk-adaptive navigation approach, a novel extension of the CVaR-BF framework that dynamically adjusts the risk level to maintain safety while ensuring trajectory feasibility. Instead of relying on a pre-specified risk parameter, our adaptive controller initializes the system with a conservative risk level (e.g., zero) and incrementally increases it until an optimization-feasible solution is achieved. In highly dynamic scenarios, where obstacles move unpredictably and rapidly, the robot requires sufficient time and space to respond and adjust its risk level. However, overly conservative strategies can limit feasible solutions in crowded environments [10, 11, 12], so an approach that maintains safety without excessively restricting the decision space is essential. To this end, we introduce the concept of a “Dynamic Zone”, where the original safety distance is adaptively expanded based on the relative position and velocity between the robot and its surrounding obstacles. The robot adjusts its trajectory before nearing obstacles, but only when needed to avoid unnecessary conservatism, while also extending the risk-adaptive range.
The main contributions of this paper are as follows:
Figure 1: Comparison of fixed vs adaptive risk levels with and without a dynamic zone.
(1) We introduce an adaptive strategy that adjusts the risk level to adopt the minimum necessary risk when navigating through obstacles, ensuring CVaR safety is guaranteed at least a pre-defined threshold while improving optimization feasibility.
(2) To increase the responsiveness to dynamic obstacles, we design a dynamic zone-based barrier function, which expands the available adjustment space for the risk level while maintaining the desired probabilistic safety guarantee.
(3) Empirical results demonstrate that our approach outperforms state-of-the-art methods in highly dynamic environments, achieving better success rates and more robustness under various of uncertain settings.
II Related Work
II-1 Risk-Aware Control under Uncertainties
Risk-aware control under uncertainties can be approached using a variety of risk measures. The approach in [1] achieves finite-time risk bounds by considering expected risk and penalizing collisions in the cost function. Similarly, [2, 13] develop probabilistic safety bounds over a finite time horizon using a discrete-time CBF condition. However, these approaches prioritize average behavior and do not provide safety guarantee at each time step. Recently, distributionally robust optimization has been designed to enhance the safety by considering a set of possible distributions to optimize decisions for worst-case scenarios. [4] reformulates the collision avoidance problem by computing safe halfspaces from obstacle sample trajectories via distributionally-robust optimization. [3] considers the reachable set of the states of an obstacle under worst-case noise. Although these methods offer robust safety guarantees, they tend to be overly conservative, which can lead to infeasible solutions in dense environments.
II-2 Dynamic Obstacle Avoidance
Geometric techniques such as Velocity Obstacles (VO) [14] and Optimal Reciprocal Collision Avoidance (ORCA) [9] utilize a velocity-obstacle framework that considers both the robot’s and obstacles’ velocities for avoidance. However, these methods typically do not account for the robot’s dynamic model. Another popular trend involves learning-based methods [15]. For instance, [16] prevents the robot from encroaching on the intended paths of other obstacles. Despite their promise, these methods often lack safety analysis and face generalization challenges when dealing with out-of-distribution data.
For safety-critical navigation, a rigorous theoretical framework and analysis are required. In [17], a coupled planning approache is proposed to model the obstacles’ motion and solve a joint optimization problem with explicit safety constraints. However, coupled methods falter when the obstacle model is inaccurate [18], and accurately modeling obstacle motion is challenging. CBFs [8] have been developed as safety filters that mitigate unsafe control inputs. Some approaches partition the solution space into convex regions using separating hyperplanes [19, 4]. For example, [4] reformulates the collision avoidance problem by computing safe halfspaces based on dynamic obstacle sample trajectories. Recent work has proposed a CBF approach combined with velocity obstacles [10, 11, 12], where the velocity obstacle principle allows the vehicle to maintain forward motion without compromising safety. However, this method tends to trigger avoidance maneuvers even when obstacle are far away, thereby leading to an conservative behavior.
III Preliminaries
III-A Discrete-Time Control Barrier Functions
Consider a robot whose motion is modeled by a discrete-time control system:
$\mathbf{x}_{k+1}=f(\mathbf{x}_{k},\mathbf{u}_{k}),$ (1)
where $\mathbf{x}_{k}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state at time step $k\in\mathbb{Z}^{+}$, and $\mathbf{u}_{k}\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, with $\mathcal{U}$ being the set of admissible controls for system (1). The function $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is assumed to be locally Lipschitz continuous with respect to $\mathbf{x}_{k}$ and $\mathbf{u}_{k}$.
According to [20], the function $h$ is a discrete-time CBF for system (1) if there exists an extended class $\mathcal{K}_{\infty}$ function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ such that there exists a control input $\mathbf{u}_{k}$ satisfying
$\Delta h(\mathbf{x}_{k},\mathbf{u}_{k})\coloneqq h(\mathbf{x}_{k+1})-h(\mathbf{x}_{k})\geq-\alpha\big{(}h(\mathbf{x}_{k})\big{)}.$ (2)
Following [20], we choose a linear function $\alpha(r)=\gamma\,r$ with $0<\gamma\leq 1$, so that the condition (2) becomes
$h(\mathbf{x}_{k+1})\geq(1-\gamma)h(\mathbf{x}_{k}),$ (3)
which guarantees that the value of $h(\mathbf{x})$ decays at an exponential rate governed by $1-\gamma$. Further, given a nominal control input $\bar{\mathbf{u}}_{k}$, a CBF-based controller can be constructed to minimally modify $\bar{\mathbf{u}}_{k}$ to guarantee safety:
$\min_{\mathbf{u}_{k}\in\mathcal{U}}\|\mathbf{u}_{k}-\bar{\mathbf{u}}_{k}\|^{2},\enspace\text{s.t.}\enspace h(\mathbf{x}_{k+1})\geq(1-\gamma)h(\mathbf{x}_{k}).$ (4)
III-B Probabilistic Constraints
Due to the inherent uncertainty in the dynamic obstacle scenarios, the safe states of the robot cannot be computed deterministically. Therefore, a probabilistic formulation of safety is required. The estimated obstacle state at time $k$, denoted $\hat{\mathbf{x}}^{o}_{k}\in\mathcal{O}\subset\mathbb{R}^{n}$, is obtained via sensor data and state estimation, and is treated as a deterministic value in our framework. However, $\mathbf{x}^{o}_{k+1}$ is modeled as a random variable, reflecting the inherent uncertainty in the predicted motion of the obstacle. Define the safe set as
$\mathcal{S}=\{\mathbf{x}_{k}\in\mathcal{X}:h(\mathbf{x}_{k},{\color[rgb]{0,0,0}\hat{\mathbf{x}}^{o}_{k})}\geq 0\}.$ (5)
$h:\mathcal{X}\times\mathcal{O}\to\mathbb{R}$ is the CBF that depends on both the system state $\mathbf{x}_{k}$ and the estimated obstacle state $\hat{\mathbf{x}}^{o}_{k}$ at the current time step $k$. Therefore, the CBF at the next time step $k+1$ considering the predicted state, which is denoted as $h_{k+1}\coloneqq h\big{(}\mathbf{x}_{k+1},\mathbf{x}^{o}_{k+1}\big{)},$ becomes a random variable due to the uncertainty in the obstacle state $\mathbf{x}^{o}_{k+1}$  [^1] For notational brevity, we omit the explicit argument of the CBF $h$ when its dependency is clear from the context..
To account for this uncertainty, we enforce safety to the system using the following probabilistic constraints:
$\mathbb{P}\bigl{(}h_{k+1}\geq 0\bigr{)}\geq 1-\beta,$ (6)
where $\beta\in(0,1)$ indicates the allowable probability of collisions. According to [21], this expression is equivalent to the Value-at-Risk (VaR) definition:
$\operatorname{VaR}_{\beta}(h_{k+1})=\sup_{\zeta\in\mathbb{R}}\{\zeta\mid\mathbb{P}(h_{k+1}\geq\zeta)\geq 1-\beta\},$ (7)
where $\zeta$ is a decision variable. It measures the $\beta$-quantile value of a random variable $h_{k+1}$. We then define the CVaR:
Definition 1 (Conditional Value-at-Risk (CVaR) [22]).
The expected loss in the $\beta$-tail of a random variable $h_{k+1}$, given the threshold VaRβ is described as:
$\operatorname{CVaR}_{\beta}(h_{k+1})\coloneqq\mathbb{E}[h_{k+1}\mid h_{k+1}\leq\operatorname{VaR}_{\beta}(h_{k+1})].$ (8)
It is shown that (8) can be solved by the following optimization problem
$\operatorname{CVaR}_{\beta}(h_{k+1})=-\inf_{\zeta\in\mathbb{R}}\mathbb{E}\left[\zeta+\frac{(-h_{k+1}-\zeta)_{+}}{\beta}\right],$ (9)
where $(\cdot)_{+}=\max\{\cdot,0\}$. Note, a value of $\beta\to 1$ corresponds to a risk-neutral case, i.e., $\operatorname{CVaR}_{\beta\to 1}(h_{k+1})=\mathbb{E}(h_{k+1})$; whereas a value of $\beta\to 0$ is a risk-averse case, i.e., $\operatorname{CVaR}_{\beta\to 0}(h_{k+1})=\operatorname{VaR}_{\beta\to 0}(h_{k+1})$ [23]. Compared with VaR, CVaR adheres to a group of axioms crucial for rational risk assessment [24].
Now, the connection between the probabilistic constraint, VaR, and CVaR constraint is:
$\operatorname{CVaR}_{\beta}(h_{k+1})\geq 0\ \Rightarrow$ (10)
$\quad\operatorname{VaR}_{\beta}(h_{k+1})\geq 0\ \Leftrightarrow\ \mathbb{P}\bigl{(}h_{k+1}\geq 0\bigr{)}\geq 1-\beta.$
III-C Discrete-Time CVaR Barrier Functions
To handle uncertainty in a risk-aware manner, we employ a dynamic coherent risk measure known as CVaR-Safety [5]. Define the cumulative CVaR values from time 00 to $k$ as the composition of per-step CVaR operators applied sequentially over the time horizon. Formally, it is defined as:
$\operatorname{CVaR}_{\beta}^{0:k}:=\operatorname{CVaR}_{\beta}^{0}\circ\operatorname{CVaR}_{\beta}^{1}\circ\dots\circ\operatorname{CVaR}_{\beta}^{k}.$
Definition 2 (CVaR-Safety [5]).
Given a safe set $\mathcal{S}$ as defined in (5) and a risk level $\beta\in(0,1)$, we call the solutions to (1), starting at $\mathbf{x}_{0}\in\mathcal{S}$, CVaR-safe if
$\operatorname{CVaR}_{\beta}^{0:k}(h_{k})\geq 0,\quad\forall\,k\geq 0.$ (11)
To enforce CVaR-safety, we utilize CVaR barrier functions for discrete-time systems.
Definition 3 (CVaR Barrier Functions [5]).
For the discrete-time system (1) and a risk level $\beta\in(0,1)$, a continuous function $h:\mathbb{R}^{n}\to\mathbb{R}$ is called a CVaR barrier function for the safe set $\mathcal{S}$ (5) if there exists a constant $\gamma\in(0,1]$ such that for each $\mathbf{x}_{k}\in\mathbb{R}^{n}$, there exist a $\mathbf{u}_{k}\in\mathbb{R}^{m}$ such that:
$\operatorname{CVaR}_{\beta}^{k}\bigl{(}h_{k+1}\bigr{)}\geq(1-\gamma)\,h_{k},\quad\forall\,\mathbf{x}_{k}\in\mathcal{X}.$ (12)
Theorem 1 ([5]).
Consider the discrete-time system in (1) and the safe set $\mathcal{S}$ as defined in (5). Let $\beta\in(0,1)$ be a given confidence level. Then, $\mathcal{S}$ is CVaR-safe if there exists a CVaR barrier function as defined above.
Proof.
The proof can be found in [5]. ∎
Figure 2: Illustration of relationship between CVaR, the risk level $\beta$, and safety. The shaded blue region represents the boundary of $\operatorname{CVaR}_{\beta}(h_{k+1})$, which includes: (a) all possible obstacle positions, (b) the portion of the distribution captured by the CVaR at the specified risk level $\beta$, and (c) the expected (mean) positions of the obstacles. Here, $p$ refers to a Probabilistic Mass Function (PMF).
Similar to the constraint function of the CBF for deterministic systems in (4), the optimization problem using the CVaR-BF constraint can be constructed as
$\min_{\mathbf{u}_{k}\in\mathcal{U}}\|\mathbf{u}_{k}-\bar{\mathbf{u}}_{k}\|^{2},\,\text{s.t.}\,\operatorname{CVaR}_{\beta}^{k}\bigl{(}h_{k+1}\bigr{)}\geq(1-\gamma)\,h_{k}.$ (13)
IV Adaptive Risk Level of CVaR-BF
IV-A Safety and Feasibility Analysis
We identify a key issue in CVaR-BF constraint: the tuning hyper-parameter, the risk level $\beta$, presents a trade-off between the safety and feasibility of the optimization problem in (13). Specifically, this risk level $\beta$, which is set by the user and held constant throughout the trajectory [5], plays a critical role in determining the robot’s behavior near the boundary of the safe set. We analyze this issue from a set-based perspective in two aspects: safety and feasibility.
Safety Analysis: The CVaR-BF constraint ensures that the probabilistic constraint in (LABEL:eq:psafe) is satisfied. As shown in Fig. 2: (1) higher $\beta$ relaxes the CVaR-BF constraint, allowing the robot to operate closer to obstacles, albeit at the expense of increased risk; (2) lower $\beta$ enforces a tighter CVaR-BF constraint, yet may cause a higher chance of infeasibility and over-conservative decisions.
Feasibility Analysis: Define the overall feasible set at each time step $k$ as the intersection of the reachable set and the CVaR-BF constraint set. Specifically, the reachable set from the current state $\mathbf{x}_{k}$ to the next state at time $k+1$ is given by
$\mathcal{R}_{k}=\left\{\mathbf{x}_{k+1}\in\mathcal{X}\,\middle|\,\exists\,\mathbf{u}_{k}\in\mathcal{U}\text{ s.t. }\mathbf{x}_{k+1}=f(\mathbf{x}_{k},\mathbf{u}_{k})\right\}.$ (14)
Subsequently, at a given risk level $\beta$, the set of admissible control inputs that satisfy the CVaR-BF constraint at time $k$ is defined as
$\mathcal{U}^{k}_{\beta}=\left\{\mathbf{u}_{k}\in\mathcal{U}\,\middle|\,\operatorname{CVaR}_{\beta}^{k}\bigl{(}h_{k+1}\bigr{)}\geq(1-\gamma)\,h_{k}\right\}.$ (15)
Consequently, the set of states for which the CVaR-BF constraint is feasible at time step $k$ is
$\mathcal{S}_{\operatorname{CVaR},k}=\left\{\mathbf{x}_{k}\in\mathcal{X}\,\middle|\,\mathcal{U}^{k}_{\beta}\neq\emptyset\right\}.$ (16)
Thus, the overall feasible set is formulated as
$\mathcal{F}_{k}=\mathcal{R}_{k}\cap\mathcal{S}_{\operatorname{CVaR},k}.$ (17)
Notably, since the CVaR is monotonically increasing with respect to $\beta$ according to its definition in (8), increasing $\beta$ relaxes the safety constraint and consequently enlarges the feasible set. This analysis reveals a trade-off: if the safety constraints are too strict (small $\beta$), the intersection $\mathcal{F}_{k}$ may be empty, leading to infeasibility; if they are too loose (large $\beta$), safety may be compromised.
IV-B Adaptive Risk Level
To resolve this conflict, we propose an adaptive tuning strategy that at each time step $k$ adjusts $\beta$ based on the robot’s perceived risk relative to the obstacles. This adaptive mechanism aims to maintain a proper balance between feasibility and safety.
Let us define the adaptive risk level at each time $k$ as
$\beta_{k}\coloneqq\min\{\beta\in(0,\beta_{u}]\mid\mathcal{U}^{k}_{\beta}\neq\emptyset\},$ (18)
where $\beta_{u}$ is the fixed risk level used in the standard CVaR formulation as in (8) and here we use it as the upper bound of the adaptive risk level. In words, we adaptively select $\beta_{k}$ as the smallest value, not exceeding $\beta_{u}$, that yields a nonempty control space.
Definition 4 (Risk Adaptive CVaR Barrier Function).
Consider the discrete-time system (1) and an adaptive risk level $\beta_{k}$ at each time step $k$ as defined in (18). A function $h:\mathbb{R}^{n}\to\mathbb{R}$ is called a Risk Adaptive CVaR Barrier Function for the safe set $\mathcal{S}$ in (5) if there exists a constant $\gamma\in(0,1]$ such that for each $\mathbf{x}_{k}\in\mathbb{R}^{n}$, there exist a $\mathbf{u}_{k}\in\mathbb{R}^{m}$ such that,
$\operatorname{CVaR}_{\beta_{k}}^{k}\left(h_{k+1}\right)\geq(1-\gamma)\,h_{k},\quad\forall\mathbf{x}_{k}\in\mathcal{X}.$ (19)
The notion of the risk adaptive CVaR barrier function allows to initialize the risk level with a conservative value (close to zero), and then incrementally increase it when necessary. A trade-off is needed only when the robot nears obstacles, while risk level remains low elsewhere to maintain a high probability of safety throughout the trajectory  [^2] We restrict the adaptive risk level to not exceed $\beta_{u}$, so that in the worst-case scenario the risk probability remains identical to that of the fixed-parameter formulation..
Theorem 2 (CVaR-Safety with Adaptive Risk Level).
Consider the discrete-time system (1) and the safe set $\mathcal{S}$ (5). Let $\beta_{u}\in(0,1)$ be a fixed upper-bound risk level, and let $\beta_{k}\in(0,\beta_{u}]$ be an adaptive risk level at time $k$ as defined in (18). Then, $\mathcal{S}$ is at least CVaR-safe with respect to the risk level $\beta_{u}$ if there exists a risk adaptive CVaR barrier function as defined in Definition 4.
Proof.
Since by construction $\beta_{k}\leq\beta_{u}$, and because CVaR is monotonic with respect to the risk level, it follows that $\operatorname{CVaR}_{\beta_{u}}^{k}(h_{k+1})\geq\operatorname{CVaR}_{\beta_{k}}^{k}(h_{k+1})$. Thus, the condition $\operatorname{CVaR}_{\beta_{k}}^{k}(h_{k+1})\geq(1-\gamma)h_{k}$ implies that $\operatorname{CVaR}_{\beta_{u}}^{k}(h_{k+1})\geq(1-\gamma)h_{k}$. Together with Theorem 1, this inequality guarantees that the solutions to (1) are CVaR-safe with a probability level that is at least as high as that ensured by a fixed risk level $\beta_{u}$. ∎
Remark 1.
Given the maximum allowable risk level $\beta_{u}$, there may still be cases where no solution exists for (18), particularly in dynamic obstacle environments.
V Dynamic Zones for CVaR Barrier Functions
We introduce the notion of a dynamic zone-based barrier function to expand the adjustment space of the $\beta_{k}$ value, while also ensuring that the given risk level $\beta_{u}$ is met.
Effective risk management is essential for robotic navigation, particularly in environments with high-speed obstacles and uncertainties that shorten reaction times and increase collision risks. Fig. 1a shows that using a fixed risk level can render the problem infeasible when the robot approaches an obstacle. In contrast, Fig. 1b employs an adaptive risk level (as described in Sec. IV-B) that starts conservatively prompting the robot to initiate a turn early, and then relaxes the safety requirements as needed. However, without incorporating a dynamic zone that provides a virtual radius (an extra buffer), this risk adjustment can ultimately compromise safety and lead to collisions. Therefore, as shown in Fig. 1c, the combination of an adaptive risk level with a dynamic zone not only facilitates early obstacle avoidance but also provides greater flexibility for risk adjustments, thereby relaxing constraints and expanding the feasible space.
V-A Dynamic Zone–Based Barrier Function
Classical CBFs typically rely on a distance-based measure that fails to account for the motion and unpredictability of obstacles. For example, one common formulation is:
$h^{\text{D}}_{k}=\|\mathbf{p}_{k}-{\color[rgb]{0,0,0}\hat{\mathbf{p}}^{o}_{k}}\|^{2}-R_{\text{safe}}^{2},$ (20)
where $R_{\text{safe}}$ represents a threshold distance capturing the minimal allowable separation, and $\mathbf{p}_{k}\in\mathbb{R}^{d}$ and ${\color[rgb]{0,0,0}\hat{\mathbf{p}}^{o}_{k}}\in\mathbb{R}^{d}$ represent the position of robot and obstacle w.r.t. a global frame (with $n=2d$ for a $d$-dimensional space), respectively. This formulation may lead to myopic behavior, causing the robot to navigate too close to obstacles and thereby increasing the risk of collision.
In contrast, functions based on velocity obstacles or collision cones [10, 11, 12] incorporate the relative motion between the robot and the obstacle. A typical candidate is:
$h^{\text{C}}_{k}=\langle\mathbf{p}^{\text{rel}}_{k},\mathbf{v}^{\text{rel}}_{k}\rangle+\|\mathbf{p}^{\text{rel}}_{k}\|\|\mathbf{v}^{\text{rel}}_{k}\|\cos\phi,$ (21)
$\mathbf{p}^{\text{rel}}_{k}=\mathbf{p}_{k}-{\color[rgb]{0,0,0}\hat{\mathbf{p}}^{o}_{k}},\quad\mathbf{v}^{\text{rel}}_{k}=\mathbf{v}_{k}-{\color[rgb]{0,0,0}\hat{\mathbf{v}}^{o}_{k}},$
where $\phi$ is the half-angle of the collision cone, defined as $\cos\phi=\sqrt{\|\mathbf{p}^{\text{rel}}_{k}\|^{2}-R_{\text{safe}}^{2}}\,/\,\|\mathbf{p}^{\text{rel}}_{k}\|$, and $\mathbf{v}_{k}\in\mathbb{R}^{d}$ and ${\color[rgb]{0,0,0}\hat{\mathbf{v}}^{o}_{k}}\in\mathbb{R}^{d}$ represent the velocity of robot and obstacle, respectively. This CBF enforces that the angle between $\mathbf{p}^{\text{rel}}_{k}$ and $\mathbf{v}^{\text{rel}}_{k}$ remains less than $\pi-\phi$, thereby ensuring that the robot is directed away from the obstacle. However, such geometric constraints can be overly conservative. In highly dynamic environments, they may force the robot to execute unnecessarily, leading to premature path diversions and, in some cases, rendering the navigation problem infeasible [12].
To address these limitations, we propose a dynamic zone–based barrier function that leverages the predicted relative state:
$h^{\text{Z}}_{k}$ $\coloneqq\|\mathbf{p}_{k}-{\color[rgb]{0,0,0}\hat{\mathbf{p}}^{o}_{k}}\|^{2}-R_{\text{safe}}^{2}\Bigl{(}1+\Delta_{k}\Bigr{)},$ (22)
$\Delta_{k}$ $=\left(-\frac{\langle\mathbf{p}^{\text{rel}}_{k},\mathbf{v}^{\text{rel}}_{k}\rangle}{\|\mathbf{p}^{\text{rel}}_{k}\|\,\|\mathbf{v}^{\text{rel}}_{k}\|}\right)_{+},$
where $(\cdot)_{+}$ denotes the nonnegative part, i.e., $\max\{0,\cdot\}$, ensuring that $\Delta_{k}\in[0,1]$  [^3] To facilitate the optimization, we employ the softplus function as an approximation of the max operator to get nonzero gradient [25].. The interpretation is as follows:
  • If $\langle\mathbf{p}^{\text{rel}}_{k},\mathbf{v}^{\text{rel}}_{k}\rangle\geq 0$Rightarrow $\Delta_{k}=0$. This implies that the robot and the obstacle are moving away from each other, thereby reducing the likelihood of a collision.
  • If $\langle\mathbf{p}^{\text{rel}}_{k},\mathbf{v}^{\text{rel}}_{k}\rangle<0$, then $\Delta_{k}>0$, indicating that the robot and the obstacle are approaching each other, thereby increasing the likelihood of a collision.
Instead of imposing a direct constraint on the relative angle which can lead to unnecessary obstacle avoidance when the robot is far away, our approach modulates the safety zone only when necessary. When the robot and obstacles are far apart, even if the safety zone radius is expanded, it does not significantly influence the robot’s behavior due to the large relative distance. Thus, this strategy prevents unnecessary avoidance of obstacles and avoids the overly conservative behavior that can result from rigid angle constraints.
V-B Probabilistic Safety Guarantee
Next, we detail how to derive a new upper bound, denoted by $\bar{\beta}_{u}$, for the adaptive risk level within the dynamic zone–based approach. The key insight is that the dynamic zone represents a larger, dynamic, yet virtual safety distance, rather than the actual physical separation between the robot and an obstacle (see Fig. 1c). We propose an analytical formulation that leverages this expanded safety zone to permit a wider range of risk level adaptation, i.e., $\beta_{k}\in(0,\bar{\beta}_{u}]$, while preserving the same probabilistic safety guarantee as that achieved with the original conservative risk level $\beta_{u}$ associated with the conventional distance-based barrier function $h^{\text{D}}_{k}$ (20).
Let $h^{\text{D}}_{k+1}$ denote the original distance‐based function with probability density function $P_{h^{\text{D}}_{k+1}}(x)$, where $x$ represents a possible realization of $h^{\text{D}}_{k+1}$. Recall that our dynamic zone‐based barrier function $h^{\text{Z}}_{k+1}$ is defined by expanding the safety radius via a factor related to $\Delta_{k}$; equivalently, we express it as a deterministic shift:
$h^{\text{Z}}_{k+1}=h^{\text{D}}_{k+1}-\Delta_{k+1}\,R_{\text{safe}}^{2}.$ (23)
Since this shift is deterministic, it does not change the shape of the distribution of $h^{\text{D}}_{k+1}$. Thus, the probability density of $h^{\text{Z}}_{k+1}$ satisfies
$P_{h^{\text{Z}}_{k+1}}(x)=P_{h^{\text{D}}_{k+1}}\Bigl{(}x+\Delta_{k+1}\,R_{\text{safe}}^{2}\Bigr{)}.$ (24)
Moreover, by the translation invariance property of CVaR [24], we have
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}=\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}-\Delta_{k+1}\,R_{\text{safe}}^{2}.$ (25)
Substituting these relations (23) and (25) into the original CVaR-BF constraint yields:
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}\geq(1-\gamma)\,h^{\text{Z}}_{k}\Leftrightarrow$ (26)
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}\geq(1-\gamma)h^{\text{D}}_{k}+\bigl{(}\Delta_{k+1}-(1-\gamma)\Delta_{k}\bigr{)}R_{\text{safe}}^{2}.$
Since CVaR is monotonic with respect to the risk level, this additional term allows us to define a new, less conservative upper bound on the risk level. One can show that
$\operatorname{CVaR}^{k}_{\bar{\beta}_{u}}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}=$ $\operatorname{CVaR}^{k}_{\beta_{u}}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}+$ (27)
$\bigl{(}\Delta_{k+1}-(1-\gamma)\Delta_{k}\bigr{)}R_{\text{safe}}^{2},$
where $\beta_{u}$ is the original risk level upper bound. Thus, by adopting the dynamic zone-based barrier function $h^{\text{Z}}_{k+1}$ and selecting the adaptive risk level according to
$\beta_{k}=\min\left\{\beta\in(0,\bar{\beta}_{u}]\,\middle|\,\mathcal{U}^{k}_{\beta}\neq\emptyset\right\},$ (28)
we ensure that the probabilistic safety guarantee is maintained at the level corresponding to the original $\beta_{u}$.
Lemma 1 (Equivalence of Probabilistic Safety Guarantee).
Given a safe set $\mathcal{S}$ (5) that is CVaR-safe under the risk adaptive CVaR-BF defined in Definition 4 with the distance-based barrier function $h^{\text{D}}$ and a fixed risk level upper bound $\beta_{u}$. Then, by adopting the dynamic zone–based barrier function $h^{\text{Z}}$ together with the newly derived upper bound $\bar{\beta}_{u}$ for the adaptive risk level, the resulting safety guarantee is equivalent to that provided by the original CVaR-BF. In other words, the safe set $\mathcal{S}$ remains CVaR-safe with the same probabilistic guarantee.
Proof.
Under the distance‐based formulation, the adaptive risk level $\beta_{k}$ is chosen from $(0,\beta_{u}]$ so that the risk adaptive CVaR-BF guarantees safety (by Theorem (2)). In the dynamic zone–based formulation, $\beta_{k}$ is selected from $(0,\bar{\beta}_{u}]$. By design, when the adaptive risk level in the dynamic zone approach reaches its upper bound $\bar{\beta}_{u}$, the resulting CVaR condition is equivalent to that obtained with the upper bound $\beta_{u}$ in the distance‐based case. Therefore, the safe set $\mathcal{S}$ remains CVaR‐safe with at least the same probabilistic safety guarantee. ∎
V-C Risk Adaptive CVaR-BF Optimization
Building on the CVaR-BF constraint in (4) and leveraging the adaptive risk level $\beta_{k}$ defined in (28), we formulate an optimization problem that ensures probabilistic safety while minimally deviating from a nominal control input. Consider the robot operates in an environment with $N$ dynamic, uncertain obstacles. The objective is to design a controller that drives a mobile robot toward its target while avoiding dynamic obstacles, thereby achieving a “reach and avoid” task with minimal collision probability.
Figure 3: Visualization of robot trajectories and associated metrics for the whole trajectory in an environment with 20 uncooperative obstacles. (a) Snapshot of the trajectory at a critical time step. (b) Risk level $\beta_{k}$ over time. (c) Control input over time. (d) CBF value over time (e) Distance to closest obstacle over time. In (a), when the robot and an obstacle approach each other with high relative velocity, the dynamic safety margin will expand, i.e., $\Delta_{k}>0$ making robot proactively avoid obstacles. As shown in (b), $\beta_{k}$ will also increase where feasible space is limited, but it remains below the upper bound $\bar{\beta}_{u}$, ensuring safety without being overly conservative. (e) shows robot maintains a safe distance from obstacles due to dynamic zone-based barrier function.
To approximate the CVaR term in the safety constraint, for each obstacle, we consider a finite set of possible realizations of the next obstacle state, $\mathcal{X}^{o}_{k+1}=\{\mathbf{x}^{o,(1)}_{k+1},\mathbf{x}^{o,(2)}_{k+1},\dots,\mathbf{x}^{o,(L)}_{k+1}\}$, each associated with a probability $p(\mathbf{x}^{o,(j)}_{k+1})$. These samples may be drawn from any predictive distribution or model that captures the uncertainty in the obstacle’s next state. For each realization, we compute the corresponding value of the random variable $h_{k+1}$, which represents the safety measure at time $k+1$. These samples are then used to approximate the expectation in the CVaR formulation in (8). Using the CVaR definition from (8) and its reformulation in (13), the following optimization problem is solved at each time $k$:
Problem 1 (Risk Adaptive CVaR-BF Optimization).
$\min_{\mathbf{u}_{k}\in\mathcal{U},\zeta_{i}\in\mathbb{R}}\;\|\mathbf{u}_{k}-\bar{\mathbf{u}}_{k}\|^{2}$
$\mathrm{s.t.}-\Bigl{(}\zeta_{i}+\frac{1}{\beta_{k}}\sum_{j=1}^{L}p_{j}\bigl{[}-h^{\text{Z}}_{i,j,{k+1}}-\zeta_{i}\bigr{]}_{+}\Bigr{)}\geq(1-\gamma_{i})\,h^{\text{Z}}_{i,k},$
$\quad\quad\quad\forall\,i\in\{1,\dots,N\},\forall\,j\in\{1,\dots,L\}.$
By introducing auxiliary variables $\eta_{j}=h^{\text{Z}}_{i,j,{k+1}}-\zeta_{i}$, we can equivalently reformulate the max operators $(\cdot)_{+}$ as linear inequalities [21]:
$\min_{\mathbf{u}_{k}\in\mathcal{U},\zeta_{i}\in\mathbb{R},\eta_{j}\in\mathbb{R}}\;\|\mathbf{u}_{k}-\bar{\mathbf{u}}_{k}\|^{2}$ (29)
[…]
$\quad\quad-\Bigl{(}\zeta_{i}+\frac{1}{\beta_{k}}\sum_{j=1}^{L}p_{j}\,\eta_{j}\Bigr{)}\geq(1-\gamma_{i})\,h^{\text{Z}}_{i,{k}},$
$\quad\quad\quad\forall\,i\in\{1,\dots,N\},\;\forall\,j=\{1,\dots,L\}.$
Remark 2.
Feasibility condition: By setting $\beta=1$, Problem 1 reduces to the deterministic CBF-based optimization with multiple constraints and input limits. Therefore, as long as the deterministic problem is feasible, Problem 1 is always guaranteed to be feasible by setting $\beta=1$, even if no feasible solution exists for any $\beta<1$.
VI Simulations
VI-A Experimental Setup
VI-A1 Implementation Details
We set $\beta_{u}=0.5$ as desired risk level and $\bar{\beta}_{u}$ can be approximately estimated by numerically solving (27). Then, to solve (28), we discretize the interval $(0,\bar{\beta}_{u}]$ into a finite set of candidate risk levels with $B$ elements, =$\mathcal{B}=\{\beta^{(1)},\beta^{(2)},\dots,\beta^{(B)}\}$, with $0<\beta^{(1)}<\beta^{(2)}<\cdots<\beta^{(B)}\leq\bar{\beta}_{u}.$ For each $\beta\in\mathcal{B}$, we assess the feasibility of the optimization problem in (LABEL:eq:opt3) using parallel processing. Specifically, at each time step $k$, the adaptive risk level $\beta_{k}$ is determined as:
$\beta_{k}=\min\{\beta\in\mathcal{B}\mid\mathcal{U}^{k}_{\beta}\neq\emptyset\},$ (30)
VI-A2 Agent Settings
[…]
We assess our method in a widely used crowd navigation simulator within a 12m $\times$ 12m space [28] (see Fig. 3), where each obstacle follows the Social Force Model (SFM) [29] with uncooperative behavior, meaning obstacles avoid collisions only with one another, not with the robot. Their maximum speeds along each axis are chosen uniformly from $\{0.3,0.6,0.9,1.2\}$m/s, and their radii are selected from $\{0.3,0.4,0.5\}$m. The robot is modeled as a double integrator, with dynamics given by
$\mathbf{p}_{k+1}$ $=\mathbf{p}_{k}+\Delta t\,\mathbf{v}_{k}+\tfrac{1}{2}\Delta t^{2}\,\mathbf{a}_{k},$ (31)
$\mathbf{v}_{k}$ $=\mathbf{v}_{k}+\Delta t\,\mathbf{a}_{k},$
[…]
(a) SR vs. number of obstacles ($\sigma$ = 0.0)
(b) SR vs. number of obstacles ($\sigma$ = 0.05)
(c) SR vs. noise level (obstacle number = 15)
Figure 4: Success rate comparisons under different noise levels and obstacle numbers.
VI-A3 Performance Metrics
Let $m_{t},m_{s},m_{f},$ and $m_{c}$ denote the total, successful, and feasible test cases, and the number of collisions, respectively, and let $\mathcal{M}$ be the set of successful cases. For each $i\in\mathcal{M}$, $L_{i}$ and $T_{i}$ denote the trajectory length and execution time, respectively. We evaluate the following performance metrics: 1) Success Rate (SR): $\frac{m_{s}}{m_{t}}$, 2) Feasibility Rate (FR): $\frac{m_{f}}{m_{t}}$, 3) Collision Rate (CR): $\frac{m_{c}}{m_{t}}$ 4) Average Trajectory Length (ATL): $\frac{1}{m_{s}}\sum\nolimits_{i\in\mathcal{M}}L_{i}$, 5) Average Trajectory Time (ATT): $\frac{1}{m_{s}}\sum\nolimits_{i\in\mathcal{M}}T_{i}$. A test case is considered feasible if, for every time step $k$ along the planned trajectory, the feasible set $\mathcal{F}_{k}$ (17) of the optimization problem in (LABEL:eq:opt3) is nonempty, i.e., $\forall k,\;\mathcal{F}_{k}\neq\emptyset$. Furthermore, a test case is deemed successful if it is feasible and the robot reaches the goal point, denoted by $\mathbf{x}_{\text{goal}}$, i.e., $\bigl{(}\forall k,\quad\mathcal{F}_{k}\neq\emptyset\bigr{)}\,\wedge\,\bigl{(}\exists\,K:\,\|\mathbf{x}_{k}-\mathbf{x}_{\text{goal}}\|<\varepsilon,\,\forall k\geq K\bigr{)}$, where $\varepsilon$ is a small positive threshold.
VI-B Benchmark Comparisons
We compare our method with the following baselines:
  1. 1.
    CVaR-BF Methods: distance-based CVaR-BF with fixed risk levels ($\beta=0.01$ or $0.99$[5] (CVaR(Dist)).
  2. 2.
    CBF Methods: collision cone-based approaches [10], including a robust control for worst-case scenarios (RCBF(Cone)) and a standard CBF-based control without explicit uncertainty handling (CBF(Cone)).
  3. 3.
    RL Methods: socially attentive reinforcement learning (CrowdNav[28] and its extension incorporating predicted obstacle intentions (CrowdNav++[16].
  4. 4.
    Geometric Methods: reciprocal velocity obstacles for collision-free motion (ORCA[9].
Note that both ORCA and the RL methods assume a holonomic model, resulting in a simpler problem setup compared to the other methods. For all experiment, we average on 120 different configurations for the obstacles to get the results.
TABLE I: Comparison results on 5 obstacle scenario with noise $\sigma=0.05$
Method SR FR CR ATL ATT
Proposed 0.917 1.000 0.058 13.036 13.092
CVaR(Dist) $\beta=0.01$ 0.008 0.008 0.000 11.492 9.600
CVaR(Dist) $\beta=0.99$ 0.575 0.683 0.108 11.571 9.238
RCBF(Cone) 0.267 0.267 0.000 11.697 8.175
CBF(Cone) 0.450 0.750 0.300 11.456 8.117
ORCA 0.417 - 0.583 11.424 6.822
CrowdNav 0.642 - 0.358 13.459 8.260
CrowdNav++ 0.833 - 0.167 14.266 8.451
Table I presents the results for the scenario with $5$ obstacles and a noise level $\sigma$ = 0.05. As we can see, our proposed method achieves the highest success rate by achieving the highest feasibility rate while keeping a low collision rate, demonstrating our adaptive CVaR-BF helps improve optimization feasibility and safe decision. In contrast, the CVaR(Dist) methods show limitations: with $\beta=0.99$, the absence of a dynamic zone results in a higher collision rate, whereas with $\beta=0.01$, the lack of adaptive risk adjustment significantly reduces the feasibility rate. The cone-based methods, RCBF(Cone) and CBF(Cone), while yielding a lower collision rate, is overly conservative and also suffer from poor feasibility. Although CrowdNav++ achieves the second highest success rate, it lacks a probabilistic safety guarantee, leading to a high collision rate.
We provide more analysis of the statistical SR in Fig. 4 with respect to different number of obstacles and noise levels. Fig. 4a shows that in deterministic environment ($\sigma=0.0$), CrowdNav++ achieves the highest success rate since it is trained in the same deterministic environment. CrowdNav lacks obstacle trajectory prediction, thus exhibiting a lower success rate. Note that all CVaR-BF based methods reduce to standard CBF methods when $\sigma=0.0$, leading to ineffective risk adjustments, therefore does not show superior performance. However, since our approach incorporates a dynamic zone, it still outperforms cone-based and distance-based barrier functions. As shown in Fig. 4b, introducing noise reduces the SR of RL methods regardless of the number of obstacles. Although CrowdNav++ considers randomized obstacle behaviors (e.g., goal change and size variation), it still experiences a significant performance drop under uncertainty. More conservative methods, e.g., RCBF(Cone) and CVaR(Dist) with $\beta=0.01$ also suffer significant degradation due to feasibility issues, resulting in the worst performance. An interesting finding is that our method performs even better in uncertain environments (Fig. 4b) than in deterministic ones (Fig. 4a). This is because in deterministic environments, our method degrades to CBF, thus the entire risk adaptation mechanism becomes void, making it less flexible. Fig. 4b and Fig. 4c further indicate that as noise level increases, the performance of all methods deteriorates. However, our proposed approach consistently achieves the highest success rate, demonstrating its robustness against uncertainty.
VI-C Ablation Studies
Table II extends the experiments presented in Table I by using the same scenario, providing further insights on the contributions of individual components in our approach.
TABLE II: Ablation study results: performance comparison of the proposed method and its variants using different CBFs and risk-level strategies.
Method Setting SR CR FR ATL ATT
Proposed 0.917 0.058 1.000 13.036 13.092
w/o dynamic zone Dist-based 0.859 0.133 1.000 13.198 13.914
Cone-based 0.548 0.067 0.615 13.529 11.793
w/o risk adaptation Fix $\beta$=0.01 0.681 0.007 0.689 13.395 13.000
Fix $\beta$=0.99 0.741 0.052 0.793 13.253 12.566
VI-C1 Effectiveness Validation of Dynamic Zone CBF
We first replace the dynamic zone-based barrier function with other counterparts. The distance-based barrier function achieves a high feasibility rate but at the sacrifice of safety. Conversely, the cone-based barrier function is excessively conservative, leading to poor overall performance. These observations underscore the importance of the dynamic zone: even with risk adaptation, other CBF methods still struggle in such highly dynamic obstacle environments.
VI-C2 Effectiveness Validation of Risk Adaptation
We further validate the effectiveness of adaptive risk levels by using fixed risk levels. A low $\beta$ value (0.01) yields a very low collision rate, but at the expense of feasibility rates. Conversely, increasing $\beta$ to 0.99 enhances the feasibility rate, yet with an increase in collisions. In contrast, our risk adaptation achieves great feasibility at the very low safety cost.
VII Conclusions
In this work, we presented a novel risk-adaptive navigation approach based on CVaR-BF that leverages a dynamic zone-based barrier function with an adjustable risk level. Our method flexibly accommodates uncertainties in obstacle models, avoiding overly conservative behavior while maintaining high feasibility and low collision rates even in crowded, dynamic environments. The proposed approach achieves the highest success rate among all baselines, especially under significant uncertainty. Future work includes quantifying the estimation error of human trajectory prediction models, developing continuous-time formulations, and validating our method in more realistic navigation scenarios.